OpenStudy (anonymous):

A projectile is fired from ground level with an initial speed of 550 m/sec and an angle of elevation of 30 degrees. Use that the acceleration due to gravity is 9.8 m/sec^2. The range I think is 564.12 ms

OpenStudy (anonymous):

Hint: How do you break down the components of velocity?

OpenStudy (anonymous):

550cos30 = 549.9770 and 550sin30 =5.0561

OpenStudy (anonymous):

Good ol' kinematics

OpenStudy (anonymous):

yep I know ...

OpenStudy (anonymous):

but this is for calculus 3 -_- so I can't use kinematics

OpenStudy (anonymous):

What are we looking for here?

OpenStudy (anonymous):

Ah lol

OpenStudy (anonymous):

range

OpenStudy (anonymous):

I got a range of 564.12m

OpenStudy (anonymous):

\[ \sin \frac{\pi}{6} = \frac{||v_y||}{550} \\ = 550 \sin \frac{\pi}{6} = ||v_y|| \\ = 275 \\ \cos \frac{\pi}{6} = \frac{||v_x||}{550} \\ = 550 \cos \frac{\pi}{6} = ||v_x|| = 476.313\]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

You're doing this in Calculus 3? Isn't this physics?

OpenStudy (anonymous):

yea it is calculus 3 LOL -_-

OpenStudy (anonymous):

I thought it was for physics as well, now he has me worried for calc 3 :x

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

my first exam in cal 3 was 72.5 =( cry

OpenStudy (anonymous):

range came out to be 26731m